OpenStudy (anonymous):

Write a quadratic equation of the form x^2 + bx + c = 0 that has the roots of 8 and 11

OpenStudy (anonymous):

roots of 8 and 11 means x - 8 = 0 and x - 11 = 0 x = 8 x = 11 so we have...(x-8)(x-11) = 0 expand and we have.. x^2-19x+88=0 hope this helped!